Start learning →

The probability of getting at least one head when a coin is tossed twice is:

A$\tfrac14$
B$\tfrac34$
C$1$
D$\tfrac12$
Answer & Solution
Correct answer: B. $\tfrac34$
P(at least one head) = 1 − P(no head) = 1 − 1/4 = 3/4.
Practise this question with Quill — AP Calculus AB →

Keep practising Unit 8B: Volumes of Revolution and Cross-Sections

Browse all Unit 8B: Volumes of Revolution and Cross-Sections questions for AP Calculus AB, or explore more Mathematics topics.

Start a Unit 8B: Volumes of Revolution and Cross-Sections lesson →